OpenStudy (anonymous):

How would you find x&y intercept of Quadratic Function (I think) of "f(x) = -3x^2 - 6x +1 Could you include detail?

OpenStudy (mimi_x3):

x-intercept: y=0 -3x^2-6x+1=0 Use the quadratic formula to find x y-intercept:x=0 y = -3(0)^2 -6(0) + 1 =